# New Permutation Decomposition Techniques For Efficient Homomorphic Permutation

### v5

@@ C.2. Diag-to-Row Transformation
For the $U^{\tau}$ applying to any $d\times d$ matrix $A$, its diagonal distribution is $[0,d^{2}-d]$ with a common difference of $d$. Such a distribution is not symmetric around the zero diagonal. In this case, we may adjust the objective of the search algorithm: let $U_{R}^{\tau}$ only contain non-zero entries at the 00-th and $r_{c}$-th diagonals for $r_{c}=d\cdot\lceil(d-1)/2\rceil$, such that the diagonal distribution of $U_{L}^{\tau}$ converges approximately to $[0,(d^{2}-d)/2]$.
@@
Appendix D Benes network-based Homomorphic Permutation Implementation
@@
  • The Benes network generated for a permutation of length $n$ typically requires a great multiplication depth of $12\log{n}-$ 1 which is unlikely to be practical. Thus, we restrict its depth to $\log{n}-1$ using the optimal level-collapsing method proposed by Halevi and Shoup (Halevi and Shoup, 2014). This depth configuration aligns with our new network construction, which is convenient for comparison. The complexity comparison in Table 2 is obtained with this setting. With depth $\log{n}-1$ being a bottom line, we further search for the best performance of Benes network-based homomorphic permutation by collapsing more levels. This leads to the results in Table 7.
@@ Appendix E Oblivious Neural Network Inference Implementation
for $10\leq i^{\prime},j^{\prime}<d_{K}=\left\lceil\frac{(n_{i}-k)}{s}\right\rceil+$ 1. In E2DM-lite we have $k=7,s=3,h=2,d_{K}=8$, and $n_{i}=28$. The encoding method for input of homomorphic convolution is similar to that of E2DM (Jiang et al., 2018). We first duplicate the set $\{I_{si^{\prime}+i,sj^{\prime}+j}\}_{0\leq i^{\prime},j^{\prime}<d_{K}}$ $h$ times to form one column of a matrix $I^{(i,j)}$, which has $b$ such columns for $b$ images to be processed simultaneously. On the other hand, the set $\{K_{i,j}^{(l)}\}_{0\leq l<h}$ is vertically concatenated and then duplicated $d_{K}^{2}\times b$ times to form a matrix $K^{(i,j)}$ with the same dimensions as $I^{(i,j)}$. Encrypting $I^{(i,j)}$ and $K^{(i,j)}$ into a ciphertext, respectively, for $0\leq i,j<k$, the homomorphic evaluation of the convolution layer can be computed as:
@@
The FC-1 Layer. In the plaintext scenario, this step computes the matrix multiplication of a $w\times hd_{K}^{2}$ weight matrix and an input matrix of size $hd_{K}^{2}\times b$ (for E2DM-lite $w=32$). In the ciphertext scenario, this is achieved through HMM across several submatrices. Specifically, when using an HMM of dimension $d$, we partition $I^{(i,j)}$ (and $K^{(i,j)}$) into $hd_{K}^{2}/d\times b/d$ submatrices, ensuring that $ct(C^{2})$ maintains the same partitioning. We also partition the weight matrix into $w/d\times hd_{K}^{2}/d$ submatrices. This allows us to utilize the parallelism of HMM for $wb/d^{2}$ groups of matrix multiplications, each involving $d\times d$ submatrices of size $hd_{K}/d$. The results are then aggregated in a group-wise order to obtain an output $ct(F)$ with $F$ of size $w\times b$.
@@
Table 9 presents benchmarks for E2DM-lite, with $\log{N}=15$ and an encoding space of length $2^{14}$. When using Jiang et al.’s HMM, with $d=32$ and $b=128$, the FC-1 layer performs 4 groups of 4 matrix multiplications of size $32\times 32$, which, with parallelism support, results in a complexity comparable to performing a single $32\times 32$ matrix multiplication. When using our HMM, we set three different parameter sets: $(b=32,d=32,k=4,d_{0}=8)$, $(b=16,d=16,k=4,d_{0}=4)$, and $(b=8,d=8,k=4,d_{0}=2)$. This transforms the FC-1 layer into performing 1, 2, and 4 groups of matrix multiplications of size $32\times 32$, $16\times 16$, and $8\times 8$, respectively. The low latency shown in the table is mainly due to the following reasons: (i) our HMM uses $k=4$-fold redundancy encoding space to provide faster computational speed. (ii) The ciphertext output from the FC-1 layer contains $2^{14}/(32b)$ duplication of the expected matrix of size $32\times b$, which is sufficient to allow the FC-2 layer to collapse into a single multiplication and a recursive summation with $\log{32}$ rotations. This enables the entire E2DM-lite computation to be completed at a lower modulus level, where each layer’s computational cost is lower than that of the HMM implementation by Jiang et al (Jiang et al., 2018).

### v7

@@ C.2. Diag-to-Row Transformation
For the $U^{\tau}$ applying to any $d\times d$ matrix $A$, its diagonal distribution is $0,d^{2}-d]$ with a common difference of $d$. Such a distribution is not symmetric around the zero diagonal. In this case, we may adjust the objective of the search algorithm: let $U_{R}^{\tau}$ only contain non-zero entries at the 0-th and $r_{c}$-th diagonals for $r_{c}=d\cdot\lceil(d-1)/2\rceil$, such that the diagonal distribution of $U_{L}^{\tau}$ converges approximately to $0,(d^{2}-d)/2]$.
@@
Appendix D Section 5 Supplemental Materials
D.1. Correctness of the Multi-group Network Construction
We prove that given any permutation $p$ of length $n$ and any $n$-dimensional input vector $\mathbf{v}$, the proposed multi-group network structure correctly outputs $p(\mathbf{v})$ with at most $\lceil\log{(\max\{r_{org}^{(i)}\})}\rceil$ network levels for $0\leq i<n$, where $r_{org}^{(i)}$ is the initial remaining rotation distance of the $i$-th entry in $\mathbf{v}$ (Recall Section 5).
Let us keep in mind that the $i$-th entry of $\mathbf{v}$ will end up being the $i+r_{org}^{(i)}$-th entry of $p(\mathbf{v})$. At any point during the traversal of $\mathbf{v}$’s entries through the network, let $r_{rem}^{(i)}$ represent the $i$-th entry’s current remaining rotation distance to its destination, initially set as $r_{org}^{(i)}$. Note that when $r_{rem}^{(i)}=0$, the entry would have reached its destination. We now prove by induction that for any $0\leq i<n$, the value $r_{rem}^{(i)}$ will be reduced to 0 after at most $\lceil\log({\max\{r_{org}^{(i)}\mid 0\leq i<n\})}\rceil$ levels.
When entries are initially passed from level 0 to $1$, we have $r_{rem}^{(i)}=r_{org}^{(i)}$ for $0\leq i<n$. At this level, rotation nodes from all groups ensure that $r_{rem}^{(i)}\leftarrow r_{rem}^{(i)}-rot$ whenever $r_{rem}^{(i)}>rot$, where $rot=2^{\lfloor\log{m}\rfloor}$ and $m=\max\{r_{org}^{(i)}\mid 0\leq i<n\}$. Consequently, by the end of level 1, the maximum remaining rotation distance across all entries will have been reduced to less than $\max\{r_{org}^{(i)}\mid 0\leq i<n\}/2$.
Assume inductively that upon entering level $k+1$, the current maximum $r_{rem}^{(i)}$ is less than $\max\{r_{org}^{(i)}\mid 0\leq i<n\}/2^{k}$. Then, as in the base case, the rotation nodes apply a subtraction of $rot=2^{\lfloor\log{m}\rfloor}$ to any $r_{rem}^{(i)}>rot$, where $m=\max\{r_{rem}^{(i)}\mid 0\leq i<n\}$. Thus, by the end of level $k+1$, the maximum $r_{rem}^{(i)}$ is again halved, falling below $\max\{r_{org}^{(i)}\mid 0\leq i<n\}/2^{k+1}$.
Therefore, after $\ell=\lceil\log{(\max\{r_{org}^{(i)}\mid 0\leq i<n\})}\rceil$ levels, we have $\max\{r_{rem}^{(i)}\mid 0\leq i<n\}<\max\{r_{org}^{(i)}\mid 0\leq i<n\}/2^{\ell}$, which equals zero.
D.2. Benes network-based Homomorphic Permutation Implementation
@@
  • The Benes network generated for a permutation of length $n$ typically requires a great multiplication depth of $2\log{n}-1$ which is unlikely to be practical. Thus, we restrict its depth to $\log{n}-1$ using the optimal level-collapsing method proposed by Halevi and Shoup (Halevi and Shoup, 2014). This depth configuration aligns with our new network construction, which is convenient for comparison. The complexity comparison in Table 2 is obtained with this setting. With depth $\log{n}-1$ being a bottom line, we further search for the best performance of Benes network-based homomorphic permutation by collapsing more levels. This leads to the results in Table 7.
@@ Appendix E Oblivious Neural Network Inference Implementation
for $0\leq i^{\prime},j^{\prime}<d_{K}=\left\lceil\frac{(n_{i}-k)}{s}\right\rceil+1$. In E2DM-lite we have $k=7,s=3,h=2,d_{K}=8$, and $n_{i}=28$. The encoding method for input of homomorphic convolution is similar to that of E2DM (Jiang et al., 2018). We first duplicate the set $\{I_{si^{\prime}+i,sj^{\prime}+j}\}_{0\leq i^{\prime},j^{\prime}<d_{K}}$ $h$ times to form one column of a matrix $I^{(i,j)}$, which has $b$ such columns for $b$ images to be processed simultaneously. On the other hand, the set $\{K_{i,j}^{(l)}\}_{0\leq l<h}$ is vertically concatenated and then duplicated $d_{K}^{2}\times b$ times to form a matrix $K^{(i,j)}$ with the same dimensions as $I^{(i,j)}$. Encrypting $I^{(i,j)}$ and $K^{(i,j)}$ into a ciphertext, respectively, for $0\leq i,j<k$, the homomorphic evaluation of the convolution layer can be computed as:
@@
The FC-1 Layer. In the plaintext scenario, this step computes the matrix multiplication of a $w\times hd_{K}^{2}$ weight matrix and an input matrix of size $hd_{K}^{2}\times b$ (for E2DM-lite $w=32$). In the ciphertext scenario, this is achieved through HMM across several submatrices. Specifically, when using an HMM of dimension $d$, we partition $I^{(i,j)}$ (and $K^{(i,j)}$) into $hd_{K}^{2}/d\times b/d$ submatrices, ensuring that $ct(C^{2})$ maintains the same partitioning. We also partition the weight matrix into $w/d\times hd_{K}^{2}/d$ submatrices. This allows us to utilize the parallelism of HMM for $wb/d^{2}$ groups of matrix multiplications, each involving $d\times d$ submatrices of size $hd_{K}^{2}/d$. The results are then aggregated in a group-wise order to obtain an output $ct(F)$ with $F$ of size $w\times b$.
@@
Table 9 presents benchmarks for E2DM-lite, with $\log{N}=15$ and an encoding space of length $2^{14}$. When using Jiang et al.’s HMM, with $d=32$ and $b=128$, the FC-1 layer performs 4 groups of 4 matrix multiplications of size $32\times 32$, which, with parallelism support, results in a complexity comparable to performing a single $32\times 32$ matrix multiplication. When using our HMM, we set three different parameter sets: $(b=32,d=32,k=4,d_{0}=8)$, $(b=16,d=16,k=4,d_{0}=4)$, and $(b=8,d=8,k=4,d_{0}=2)$. This transforms the FC-1 layer into performing 1, 2, and 4 groups of matrix multiplications of size $32\times 32$, $16\times 16$, and $8\times 8$, respectively. The low latency shown in the table is mainly due to the following reasons: (i) our HMM uses $k=4$-fold redundancy encoding space to provide faster computational speed. (ii) The ciphertext output from the FC-1 layer contains $2^{14}/(32b)$ duplication of the expected matrix of size $32\times b$, which is sufficient to allow the FC-2 layer to collapse into a single multiplication and a recursive summation with $\log{32}$ rotations. This enables the entire E2DM-lite computation to be completed at a lower modulus level, where each layer’s computational cost is lower than that of the HMM implementation by Jiang et al. (Jiang et al., 2018).